kenzie126:

Solve for x: |x| − 8 = −5 A- x = −13 and x = −3 B- x = 3 and x = −3 C- x = 3 and x = 13 D- No solution

supie:

\[|x|-8=-5\] +8 (to both sides) which looks like this\[|x|-8+8=-5+8\] so \[|x|=3\] So what do you think the answer is?